# Technical Report: Enhancing LLM Reasoning with Reward-guided Tree Search

### v1

@@ Abstract
Recently, test-time scaling has garnered significant attention from the research community, largely due to the substantial advancements of the o1 model released by OpenAI. By allocating more computational resources during the inference phase, large language models (LLMs) can extensively explore the solution space by generating more thought tokens or diverse solutions, thereby producing more accurate responses. However, developing an o1-like reasoning approach is challenging, and researchers have been making various attempts to advance this open area of research. In this paper, we present a preliminary exploration into enhancing the reasoning abilities of LLMs through reward-guided tree search algorithms. This framework is implemented by integrating the policy model, reward model, and search algorithm. It is primarily constructed around a tree search algorithm, where the policy model navigates a dynamically expanding tree guided by a specially trained reward model. We thoroughly explore various design considerations necessary for implementing this framework and provide a detailed report of the technical aspects. To assess the effectiveness of our approach, we focus on mathematical reasoning tasks and conduct extensive evaluations on four challenging datasets, significantly enhancing the reasoning abilities of LLMs.
@@ 1 Introduction
Implementing a search-based reasoning framework, as demonstrated in prior studies, often entails a multitude of design considerations or technical details, As a result, reproducing an o1-like system is not trivial, even when performance is not the primary objective. A comprehensive report detailing the exploration of various implementation aspects would be useful in guiding future research in this area. In fact, several studies and projects zhang2024llamaberry ; qin2024o1replicationjourneystrategic have been dedicated to serving this purpose. In parallel with these works, we present a technical report documenting our team’s preliminary exploration of a reward-guided tree search framework for improving LLM reasoning. Our framework is primarily constructed around a tree search algorithm, where the policy model navigates a dynamically expanding tree, guided by a specially trained reward model. We thoroughly explore the various design considerations necessary to implement this framework and report the technical details. To summarize, we list the major attempts at each aspect below:
@@ 2 Method
In this work, we focus on the mathematical domain, specifically addressing mathematical problems described in the text. We implement a reward-guided tree search framework designed to enhance the reasoning capabilities of LLMs. This framework consists of three main components: the policy model, the reward model, and the search algorithm. Within our framework, the policy model generates new reasoning steps based on a partial solution prefix along a given path in the search tree. The search algorithm constructs the search tree to structure the reasoning process, while the reward model provides feedback signals to guide the policy model’s actions and the search process. This approach allows the policy model to explore a broader reasoning space in a more informed manner, increasing the likelihood of finding the correct answer. Overall, it trades test time for improved accuracy. In the following sections, we will provide detailed descriptions of the implementation of these three key components.
Figure 1: An overview of the implemented reasoning framework in this work.
@@ 3.2.1 Overall Performance Comparison
As the main experiments, we compare four methods based on the same backbone model: zero-shot CoT (CoT) with fine-tuned policy model, best-of-$N$ chosen by the reward model (BoN, $N=100$), and our tree search based reasoning framework (T-Search). Specially, We utilize LLaMA-3.1-8B-Instruct LLM as the backbone model for policy model and reward model. And we utilize the same prompt for all methods as described in Section 2.1.1.
We first present the overall performance comparison of the four methods on the selected evaluation benchmarks. As shown in Table 1, the CoT strategy can improve the performance of the original chat model to some extent, as these test tasks require enhanced step-by-step reasoning capabilities. Among these methods, search-based approaches (i.e., BoN and T-Search) deliver more superior performance because they effectively expand the solution space for challenging reasoning tasks. Finally, our framework achieves the best performance among all methods, enhancing the reference baseline by 46.9%, 7.3%, 91.6%, and 31.4% on the four evaluation benchmarks, respectively. These results indicate that our framework effectively enhances the reasoning capabilities of LLMs on complex mathematical tasks.
@@ 3.3.3 Performance Improvement on Self-consistency
To further evaluate the effectiveness of our training method for the policy model, we conduct self-consistency experiments to assess how an improved policy model contributes to model performance using the self-consistency (SC) strategy. We select the SC strategy because we empirically find that it performs very well when employing a sufficient number of rollouts. Specifically, we compare the policy model with and without (i.e., the original model) optimization using our training method, while varying the sampling budget (i.e., the number of rollouts). Additionally, we include the performance of the overall reasoning framework (i.e., T-Search in Table 1) as a reference. From Figure 3, we can see that our training method significantly enhances SC performance by improving the underlying policy model, highlighting its effectiveness. Moreover, SC with the enhanced policy model acts as a very strong baseline, achieving high performance yet still underperforming compared to our entire reasoning framework.
@@ 4 Conclusion
In this paper, we presented a technical report documenting our preliminary exploration of a search-enhanced reasoning framework for LLMs. In implementation, our framework consisted of three main components: policy model, reward model, and search algorithm. We carefully discussed various key design considerations and conducted extensive explorations of various implementation details. Preliminary results have demonstrated the effectiveness of the implemented framework on the mathematical reasoning tasks.

### v4

# Technical Report on Slow Thinking with LLMs: I

@@ Abstract
Recently, test-time scaling has garnered significant attention from the research community, largely due to the substantial advancements of the o1 model released by OpenAI. By allocating more computational resources during the inference phase, large language models (LLMs) can extensively explore the solution space by generating more thought tokens or diverse solutions, thereby producing more accurate responses. However, develop an o1-like reasoning approach is challenging, and researchers have been making various attempts to advance this open area of research. In this paper, we present a preliminary exploration into enhancing the reasoning abilities of LLMs through reward-guided tree search algorithms. This framework is implemented by integrating the policy model, reward model, and search algorithm. It is primarily constructed around a tree search algorithm, where the policy model navigates a dynamically expanding tree guided by a specially trained reward model. The implemented framework is denoted as STILL-1 (Slow Thinking with LLMs), marking the first model developed by our project, “Slow Thinking with LLMs”. We thoroughly explore various design considerations necessary for implementing this framework and provide a detailed report of the technical aspects. To assess the effectiveness of our approach, we focus on mathematical reasoning tasks and conduct extensive evaluations on four challenging datasets, significantly enhancing the reasoning abilities of LLMs.
@@ 1 Introduction
Implementing a search-based reasoning framework, as demonstrated in prior studies, often entails a multitude of design considerations or technical details, As a result, reproducing an o1-like system is not trivial, even when performance is not the primary objective. A comprehensive report detailing the exploration of various implementation aspects would be useful in guiding future research in this area. In fact, several studies and projects zhang2024llamaberry ; qin2024o1replicationjourneystrategic have been dedicated to serving this purpose. In parallel with these works, we present a technical report documenting our team’s preliminary exploration of a reward-guided tree search framework for improving LLM reasoning. Our framework is primarily constructed around a tree search algorithm, where the policy model navigates a dynamically expanding tree, guided by a specially trained reward model. We denote the implemented framework by STILL-1 (Slow Thinking LLMs), which is the first model developed by our project, “Slow Thinking with LLMs”. We thoroughly explore the various design considerations necessary to implement this framework and report the technical details. To summarize, we list the major attempts at each aspect below:
@@ 2 Method
In this work, we focus on the mathematical domain, specifically addressing mathematical problems described in text. We implement a reward-guided tree search framework, STILL-1, designed to enhance the reasoning capabilities of LLMs. This framework consists of three main components: the policy model, the reward model, and the search algorithm. Within our framework, the policy model generates new reasoning steps based on a partial solution prefix along a given path in the search tree. The search algorithm constructs the search tree to structure the reasoning process, while the reward model provides feedback signals to guide the policy model’s actions and the search process. This approach allows the policy model to explore a broader reasoning space in a more informed manner, increasing the likelihood of finding the correct answer. Overall, it trades test time for improved accuracy. In the following sections, we will provide detailed descriptions of the implementation of these three key components.
Figure 1: An overview of our reasoning framework STILL-1.
@@ 3.2.1 Overall Performance Comparison
As the main experiments, we compare four methods based on the same backbone model: zero-shot CoT (CoT) with fine-tuned policy model, best-of-$N$ chosen by the reward model (BoN, $N=100$), and our tree search based reasoning framework (STILL-1). Specially, We utilize Llama-3.1-8B-Instruct LLM as the backbone model for policy model and reward model. And we utilize the same prompt for all methods as described in Section 2.1.1.
We first present the overall performance comparison of the four methods on the selected evaluation benchmarks. As shown in Table 1, the CoT strategy can improve the performance of the original chat model to some extent, as these test tasks require enhanced step-by-step reasoning capabilities. Among these methods, search-based approaches (i.e., BoN and SKILL-1) deliver more superior performance because they effectively expand the solution space for challenging reasoning tasks. Finally, our framework achieves the best performance among all methods, enhancing the reference baseline by 46.9%, 7.3%, 91.6%, and 31.4% on the four evaluation benchmarks, respectively. These results indicate that our framework effectively enhances the reasoning capabilities of LLMs on complex mathematical tasks.
@@ 3.3.3 Performance Improvement on Self-consistency
To further evaluate the effectiveness of our training method for the policy model, we conduct self-consistency experiments to assess how an improved policy model contributes to model performance using the self-consistency (SC) strategy. We select the SC strategy because we empirically find that it performs very well when employing a sufficient number of rollouts. Specifically, we compare the policy model with and without (i.e., the original model) optimization using our training method, while varying the sampling budget (i.e., the number of rollouts). Additionally, we include the performance of the overall reasoning framework (i.e., STILL-1 in Table 1) as a reference. From Figure 3, we can see that our training method significantly enhances SC performance by improving the underlying policy model, highlighting its effectiveness. Moreover, SC with the enhanced policy model acts as a very strong baseline, achieving high performance yet still underperforming compared to our entire reasoning framework.
@@ 4 Conclusion
In this paper, we presented a technical report documenting our preliminary exploration of a search-enhanced reasoning framework STILL-1 for LLMs. In implementation, our framework consisted of three main components: policy model, reward model, and search algorithm. We carefully discussed various key design considerations and conducted extensive explorations of various implementation details. Preliminary results have demonstrated the effectiveness of the implemented framework on the mathematical reasoning tasks.